# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ C.3 Error Bound for Score Estimation
By combining the approximation and statistical errors, we obtain
$\leavevmode\nobreak\ \mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)$
$=$ $\leavevmode\nobreak\ \widetilde{\mathcal{O}}\left(\frac{1}{n}\cdot\frac{\xi^{d}\epsilon^{-(d+1)}}{(1-T)^{5}}+\frac{1}{\sqrt{m}}\cdot\frac{\xi^{\frac{d^{*}}{2}}\epsilon^{-\frac{d^{*}+1}{2}}}{(1-T)^{\frac{7}{2}}}+\epsilon^{2}\right)$
$=$ $\leavevmode\nobreak\ \widetilde{\mathcal{O}}\left(\frac{\xi^{d^{*}}}{(1-T)^{5}}\left(\frac{\epsilon^{-(d^{*}+1)}}{n}+\frac{\epsilon^{-\frac{d^{*}+1}{2}}}{\sqrt{m}}+\epsilon^{2}\right)\right).$
By setting $\epsilon=n^{-\frac{1}{d^{*}+3}}$, it holds that
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)$
$=\widetilde{\mathcal{O}}\left(\frac{\xi^{d^{*}}}{(1-T)^{5}}\left(n^{-\frac{2}{d^{*}+3}}+n^{\frac{d^{*}+1}{2(d^{*}+3)}}m^{-\frac{1}{2}}\right)\right).$
C.4 Sampling Error
Proof of Theorem 4.7.
Let
$R_{t}=\mathbb{E}{\|{\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}}\|^{2}}.$
Deriving $R_{t}$ with respect to time $t$, we have
$\frac{\mathrm{d}R_{t}}{\mathrm{d}t}$ $=2\sigma^{2}\mathbb{E}\left[\langle\nabla\log{q(\sqrt{1-t}\sigma,\mathbf{x}_{t})}-\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t}),\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}\rangle\right]$
$=2\sigma^{2}\mathbb{E}\left[\langle\nabla\log{q(\sqrt{1-t}\sigma,\mathbf{x}_{t})}-\widehat{\mathbf{s}}(t,\mathbf{x}_{t}),\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}\rangle\right]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +2\sigma^{2}\mathbb{E}\left[\langle\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t}),\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}\rangle\right].$
Using the inequality $2\langle\mathbf{a},\mathbf{b}\rangle\leq 2\|{\mathbf{a}}\|\|{\mathbf{b}}\|\leq\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}$, we obtain
$2\langle\nabla\log{q(\sqrt{1-t}\sigma,\mathbf{x}_{t})}-\widehat{\mathbf{s}}(t,\mathbf{x}_{t}),\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}\rangle$
$\leq\|\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})\|^{2}+\|{\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}}\|^{2},$
and
$2\langle\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t}),\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}\rangle$
$\leq 2\|{\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t})}\|\|{\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}}\|$
$\leq 2\sqrt{d^{*}}{\gamma_{1}}\|{\mathbf{x}_{t}-\widehat{\mathbf{x}}_{t}}\|^{2},$
where the last inequality follows from (11) with $\gamma_{1}=10d^{*}\xi$. Thus,
@@ C.4 Sampling Error
$\frac{\mathrm{d}R_{t}}{\mathrm{d}t}$ $\leq\sigma^{2}\left[(1+2\sqrt{d^{*}}\gamma_{1})R_{t}+\mathbb{E}\|{\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})}\|^{2}\right]$
$=\sigma^{2}\left[(1+2\sqrt{d^{*}}\gamma_{1})R_{t}+\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathbb{E}_{\mathbf{x}_{t}}\|{\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})}\|^{2}\right)\right].$
By Grönwall’s inequality, we have
$e^{-\sigma^{2}(1+2\sqrt{d^{*}}\gamma_{1})T}R_{T}-R_{0}$
@@
$\leq\sigma^{2}\int_{0}^{T}e^{-\sigma^{2}(1+2\sqrt{d^{*}}\gamma_{1})t}\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathbb{E}_{\mathbf{x}_{t}}\|{\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})}\|^{2}\right)\mathrm{d}t$
@@
$\leq T\sigma^{2}\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|{\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})}\|^{2}\mathrm{d}t\right).$
Since $R_{0}=0$, we finally obtain
@@
$R_{T}\leq T\sigma^{2}e^{\sigma^{2}(1+2\sqrt{d^{*}}\gamma_{1})T}\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|{\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})-\widehat{\mathbf{s}}(t,\mathbf{x}_{t})}\|^{2}\mathrm{d}t\right),$
which can be expressed as
$R_{T}=\widetilde{\mathcal{O}}\left(\frac{e^{2\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\xi^{d^{*}}}{(1-T)^{5}}\left(n^{-\frac{2}{d^{*}+3}}+n^{\frac{d^{*}+1}{2(d^{*}+3)}}m^{-\frac{1}{2}}\right)\right).$
Then, Theorem 4.7 holds by the inequality $(\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\pi_{T},\widehat{\pi}_{T})])^{2}\leq\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W^{2}_{2}(\pi_{T},\widehat{\pi}_{T})]\leq R_{T}$.
Appendix D Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\widetilde{\pi}_{T})]$
Proof of Theorem 4.8.
Let
$L_{t}=\mathbb{E}{\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2}}.$
Deriving $L_{t}$ with respect to time $t$, we have
$\frac{\mathrm{d}L_{t}}{\mathrm{d}t}$ $=2\sigma^{2}\mathbb{E}\left[\langle\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t_{k}}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\right]$
$=2\sigma^{2}\mathbb{E}\left[\langle\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t,\widetilde{\mathbf{x}}_{t}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\right]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +2\sigma^{2}\mathbb{E}\left[\langle\widehat{\mathbf{s}}(t,\widetilde{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\right]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +2\sigma^{2}\mathbb{E}\left[\langle\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t_{k}}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\right].$
Estimating each term on the right hand side of the above equality separately, we obtain
$2\langle\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t,\widetilde{\mathbf{x}}_{t}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\leq{2\gamma_{1}}\sqrt{d^{*}}\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2},$
$2\langle\widehat{\mathbf{s}}(t,\widetilde{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle\leq{d^{*}\gamma_{2}^{2}(t-t_{k})^{2}}+\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2},$
and
$2\langle\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t})-\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t_{k}}),\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}\rangle$
$\leq{d^{*}\gamma_{1}^{2}}\|{\widetilde{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t_{k}}}\|^{2}+\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2}$
$\leq{d^{*}\gamma_{1}^{2}(t-t_{k})^{2}}\|{\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t_{k}})}\|^{2}+\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2}$
$\leq{d^{*}\gamma_{1}^{2}}(t-t_{k})^{2}K^{2}+\|{\widehat{\mathbf{x}}_{t}-\widetilde{\mathbf{x}}_{t}}\|^{2},$
where $\gamma_{2}=10\tau(r)=\mathcal{O}\left(\frac{10\left(d^{*}\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{3}{2}}}{(1-T)^{3}}\right)$.
Combining the above inequalities, we have
$\frac{\mathrm{d}L_{t}}{\mathrm{d}t}\leq\sigma^{2}\left[(2\sqrt{d^{*}}\gamma_{1}+2)L_{t}+d^{*}(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2})(t-t_{k})^{2}\right].$
By Grönwall’s inequality, we have
$e^{-\sigma^{2}(2\sqrt{d^{*}}\gamma_{1}+2)t_{k+1}}L_{t_{k+1}}-e^{-\sigma^{2}(2\sqrt{d^{*}}\gamma_{1}+2)t_{k}}L_{t_{k}}$
$\leq\int_{t_{k}}^{t_{k+1}}e^{-\sigma^{2}(2\sqrt{d^{*}}\gamma_{1}+2)t}\sigma^{2}d^{*}(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2})(t-t_{k})^{2}\mathrm{d}t$
$\leq\int_{t_{k}}^{t_{k+1}}\sigma^{2}d^{*}(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2})(t-t_{k})^{2}\mathrm{d}t$
$\leq\frac{1}{3}\sigma^{2}d^{*}(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2})(t_{k+1}-t_{k})^{3}.$
Since $L_{0}=0$, we finally obtain
$L_{T}\leq\frac{1}{3}\sigma^{2}d^{*}(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2})e^{\sigma^{2}(2\sqrt{d^{*}}\gamma_{1}+2)}\sum_{k=0}^{N-1}(t_{k+1}-t_{k})^{3},$
that is,
$L_{T}=\widetilde{\mathcal{O}}\left(\left(\gamma_{1}^{2}K^{2}+\gamma_{2}^{2}\right)e^{2\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\sum_{k=0}^{N-1}(t_{k+1}-t_{k})^{3}\right).$
Then, Theorem 4.8 holds by the inequality $(\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\widetilde{\pi}_{T})])^{2}\leq\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}^{2}(\widehat{\pi}_{T},\widetilde{\pi}_{T})]\leq L_{T}$.
Appendix E Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$
Proof of Lemma 4.9.
Consider the process
where $\bar{\mathbf{w}}_{t}$ is a standard Brownian motion. Then $\bar{\mathbf{x}}_{t}\sim q(\sqrt{t}\sigma,\mathbf{x})$ also satisfies the reverse SDE
$\mathrm{d}\mathbf{x}_{t}=\sigma^{2}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t}.$
[…]
@@ Appendix E Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$
$\mathbb{E}\left[\|\mathbf{x}_{1}-\mathbf{x}_{T}\|^{2}|\mathcal{Y}\right]=\mathbb{E}\left[\|\bar{\mathbf{x}}_{0}-\bar{\mathbf{x}}_{1-T}\|^{2}|\mathcal{Y}\right]=\mathbb{E}\|\sigma\bar{\mathbf{w}}_{1-T}\|^{2}=\sigma^{2}d^{*}(1-T).$
Therefore,
@@
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]\leq\left(\mathbb{E}_{\mathcal{Y}}\mathbb{E}\left[\|\mathbf{x}_{1}-\mathbf{x}_{T}\|^{2}|\mathcal{Y}\right]\right)^{\frac{1}{2}}\leq\sigma\sqrt{d^{*}(1-T)}.$
Appendix F Oracle Inequality
Proof of Theorem 4.10.
By Assumption 4.3, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},{\widehat{p}}^{*}_{data})]$
$=\widetilde{\mathcal{O}}\left(\sqrt{1-T}+(\gamma_{1}K+\gamma_{2})e^{\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\sqrt{\sum_{k=0}^{N-1}(t_{k+1}-t_{k})^{3}}+\frac{e^{\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\xi^{\frac{d^{*}}{2}}}{(1-T)^{\frac{5}{2}}}(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}})\right).$
By the choice of neural networks, we have $\gamma_{1}=\mathcal{O}\left(\frac{10{d^{*}}^{3/2}}{\sigma^{4}(1-T)^{2}}\right)$. Let $\mathop{\mathrm{max}}_{k=0,1,\cdots,N-1}|t_{k+1}-t_{k}|=\mathcal{O}\left(n^{-\frac{2}{d^{*}+3}}+n^{\frac{d^{*}+1}{2(d^{*}+3)}}m^{-\frac{1}{2}}\right)$ and $T=1-(\log{n})^{-\frac{1}{4}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left((\log{n})^{-\frac{1}{8}}+e^{\frac{10{d^{*}}^{5/2}}{\sigma^{2}}\sqrt{\log{n}}}\left(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}}\right)\right).$
Proof of Theorem 4.11.
By Assumption 4.3 and Assumption 4.4, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},{\widehat{p}}^{*}_{data})]$
$=\widetilde{\mathcal{O}}\left(\sqrt{1-T}+\left(K+\gamma_{2}\right)\sqrt{\sum_{k=0}^{N-1}(t_{k+1}-t_{k})^{3}}+\frac{1}{(1-T)^{\frac{5}{2}}}\left(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}}\right)\right).$
By setting $\mathop{\text{max}}_{k=0,1,\cdots,N-1}|t_{k+1}-t_{k}|=\mathcal{O}\left(n^{-\frac{7}{3(d^{*}+3)}}\right)$ and $T=1-n^{-\frac{1}{3(d^{*}+3)}}$, we have

### v2

@@ D.1 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]$
For any $t\in[0,T]$, $\mathbf{x}_{t}\sim q_{1-t}(\mathbf{x})$ and $\|\widehat{\mathbf{s}}(1-t,\mathbf{x})\|$, $\|\nabla\log q_{1-t}(\mathbf{x})\|\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{1-t}}\lesssim\frac{\sqrt{\log n}}{\sqrt{1-T}}$. Therefore, it holds that
@@
$\mathbb{E}_{\mathbb{P}}\left[\exp\left(\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\frac{\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}}{2\sigma^{2}}\mathrm{d}t\right)\right]$
$\lesssim\exp\left(\mathcal{O}\left(\frac{T\log n}{1-T}\right)\right)$
$\lesssim\exp\left(\mathcal{O}\left(n^{\frac{2\beta}{d^{*}+2\beta}}\log n\right)\right)<+\infty.$
We know that condition (24) is satisfied. Therefore, by Lemma D.1 and inequality $\mathrm{TV}(\widetilde{\pi}_{T},\pi_{T})=\mathrm{TV}(\pi_{T},\widetilde{\pi}_{T})\lesssim\sqrt{\mathrm{KL}(\pi_{T},\widetilde{\pi}_{T})}$, we obtain
@@
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}^{2}(\widetilde{\pi}_{T},\pi_{T})]$ $\lesssim\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\mathbb{E}_{\mathbb{P}}\left(\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)\right]$
@@
$\lesssim\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right].$
Moreover, the term $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}^{2}(\widetilde{\pi}_{T},\pi_{T})]$ can be bounded as
@@
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[\mathrm{TV}^{2}(\widetilde{\pi}_{T},\pi_{T})]$ $\lesssim\underbrace{\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})\|^{2}\mathrm{d}t\right]}_{\mathrm{(I)}}$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\underbrace{\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right]}_{\mathrm{(II)}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\underbrace{\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right]}_{\mathrm{(III)}}.$
We bound the above three terms separately. The first term yields that
@@
$\mathrm{(I)}=\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\mathbb{E}_{\mathbb{P}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})\|^{2}(t_{i+1}-t_{i})\right].$
Therefore, for any $\epsilon>0$, there exists $\Delta_{\epsilon}>0$ such that if $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})\leq\Delta_{\epsilon}$, then it holds that